OpenStudy (anonymous):

I will give medal and become fan The distance you travel varies directly with the amount of gas you use. You can travel 300 miles on 12 gallons of gas. Question: a. You want to know how much gas it will take to for your entire road trip that is 750 miles. Write a proportion to solve this equation. b. How many gallons will it take? c. If gas is $3.75 a gallon, how much will it cost in gas to go on your road trip? d. What is the cost per mile?

OpenStudy (anonymous):

can you guys help

OpenStudy (anonymous):

@math_man21

OpenStudy (welshfella):

the proportion is 12 / 300 = x / 750 where x is the amount of gas you will use

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@priencess_j

OpenStudy (welshfella):

So to answer b you need to solve for x.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what know

OpenStudy (welshfella):

can you solve 12 / 300 = x /750 ?

OpenStudy (anonymous):

i dont know what the formula is

OpenStudy (anonymous):

@math_man21 can you help

OpenStudy (retireed):